OpenStudy (itsbribro):

Simplify. -7 + 2(x - 3) A. x - 10 B. x - 13 C. 2x - 10 D. 2x - 13

OpenStudy (michele_laino):

please, I can write: \[-7+(2*x)+(-2*3)=-7+2x-6=...\] please complete

OpenStudy (itsbribro):

D or B?

OpenStudy (michele_laino):

B no, because we have 2x., whereas in B there is x only

OpenStudy (itsbribro):

wait yea d

OpenStudy (michele_laino):

ok! that's right!

OpenStudy (anonymous):

-7+2(x-3) distribute -7 + 2x-6 add similar terms 2x-13; D = ans